# Next-Generation Database Interfaces: A Survey of LLM-based Text-to-SQL

### v6

@@ IV-A3 $\mathbf{C}_{2}$-Prompt Optimization
1. Advanced few-shot sampling: This strategy optimizes example selection through semantic similarity, diversity, or domain alignment. DESEM [63] and retrieval-augmented frameworks [66, 71, 58] de-semanticize questions to retrieve examples with matching intent skeletons or vector embeddings. DAIL-SQL [8] balances quality and quantity by masking domain-specific tokens and ranking examples based on the Euclidean distance between questions and SQL queries, while SAFE-SQL [73] combines skeleton masking with self-augmentation. ODIS [67] hybridizes out-of-domain and synthetic in-domain examples for robustness, and FUSED [69] enhances diversity through iterative demonstration synthesis. Dynamic selection is explored in ACT-SQL [68] and MCS-SQL [55], which adaptively retrieve examples based on defined similarity scores or masked similarity scores.
2. Schema augmentation: These methods refine schema representations to reduce redundancy and improve relevance. C3 [28] distills question-relevant schemas by removing unrelated tables or columns, while RSL-SQL [59] simplifies schemas using NL descriptions. ReFoRCE [61] employs table compression and column exploration to manage long-context schema complexity. Gen-SQL [60] bridges NL-SQL gaps by generating pseudo-schemas that align with user questions. QDecomp [47] and SD+SA+Voting [65] further augment prompts with schema-aware semantic and structural hints. For more complex databases, LinkAlign [62] introduces a schema-based question rewriting strategy, combined with a retrieve-then-filter schema selection mechanism for accurate linking.
3. External knowledge integration: This group enhances prompts with domain-specific knowledge or feedback mechanisms. Knowledge-to-SQL [70] trains a DELLM to inject external knowledge about question-database relationships, while ICRL [72] integrates reinforcement learning with schema-graph-derived knowledge bases. MAC-SQL [48] and DIN-SQL [7] implicitly utilize decomposition patterns as procedural knowledge to guide generation. C3 [28] and SAFE-SQL [73] incorporate calibration bias and self-augmented feedback loops, respectively, to iteratively refine outputs.
@@ IV-A4 $\mathbf{C}_{3}$-Reasoning Enhancement
1. Chain-of-thought (CoT)-based reasoning: CoT-based methods guide LLMs in generating intermediate reasoning steps before predicting SQLs. While standard CoT prompts (e.g., “Let’s think step by step”) show limited effectiveness in text-to-SQL [8, 1], adaptations help bridge this gap. ACT-SQL [68] automates CoT example generation by linking question slices to database columns via similarity matching. QDecomp [47] restructures SQL clauses into logical steps ordered by execution flow, reducing error propagation compared to generic CoT. CHASE-SQL [77] extends this with multi-path reasoning for candidate generation using “divide and conquer” CoT strategy.
2. Consistency-driven reasoning: These methods improve reliability through candidate diversity and agreement metrics before execution. Self-consistency [126] is employed in frameworks like C3 [28], DAIL-SQL [8], and SuperSQL [57], where majority voting selects the most frequent SQL candidate based on syntax or structural similarity. SD+SA+Voting [65] refines this by filtering candidates using schema-aware rules (e.g., valid joins, clause ordering) before voting. Unlike execution-dependent methods, these strategies prioritize logical consistency and syntactic validity over runtime feedback, making them lightweight but less robust to semantic errors.
3. Tool-augmented and agentic reasoning: Hybrid techniques combine SQL generation with external computation or tools. SQL-CRAFT [75] employs Program-of-Thoughts [127] (PoT) prompting, requiring LLMs to generate Python code alongside SQL for arithmetic-heavy queries, improving performance on datasets like BIRD [1]. FUXI [76] integrates dedicated tools for schema linking and syntax checking, dynamically invoking them during generation. These approaches differ from pure prompting strategies by explicitly bridging LLM reasoning with external symbolic systems. Also, recent frameworks such as ReFoRCE [61] and Spider-Agent [42] adopt step-by-step agentic reasoning, where LLMs act as specialized agents that iteratively plan actions to handle complex database scenarios.
@@ IV-A5 $\mathbf{C}_{4}$-Execution Refinement
1. Feedback-driven regeneration: These methods iteratively refine SQL queries using execution errors or intermediate results. Self-Debugging [80] enables LLMs to autonomously correct errors via natural language explanations of failed executions, while DESEM [63] uses error-type-specific prompts with termination criteria to prevent infinite loops. Multi-agent frameworks like MAC-SQL [48] and CHASE-SQL [77] utilize agents (e.g., Refiner, Query Fixer) to diagnose SQLite errors and regenerate SQL fixes. SQL-CRAFT [75] introduces adaptive control to balance correction depth, preventing over-refinement. Schema-focused methods like RSL-SQL [59] and DIN-SQL [7] use generic or gentle prompts to help LLMs identify schema misalignments during regeneration. Knowledge-to-SQL [70] optimizes its DELLM model through preference learning from execution feedback, while FUXI [76] integrates error feedback into tool-based reasoning. E-SQL [81] enhances questions with execution-aware candidate predictions. Agentic frameworks [42] also incorporate feedback-driven regeneration, where specialized agents iteratively analyze execution feedback to plan corrective actions to refine SQLs.
2. Execution-guided selection: This strategy leverages database execution results to select or rank candidates. ReFoRCE [61] and PET-SQL [52] prune invalid SQL queries that fail execution, then apply bias-aware voting (e.g., difficulty-weighted or efficiency-prioritized as in MCS-SQL [55]). Probabilistic verification distinguishes methods: MRC-EXEC [78] minimizes Bayes risk by comparing execution outputs to expected results, and LEVER [79] trains a verifier to predict correctness likelihoods from execution traces. Hybrid approaches like the retrieval-augmented framework [66] use execution feedback for iterative SQL revision through semantic gap analysis, bridging generation and selection.
@@ IV-B Fine-tuning
Enhanced Architecture CLLMs [82] DeepSeek* [11] Mar-2024 ICML’24
Pre-training CodeS [24] StarCoder [11, 1] Feb-2024 SIGMOD’24
Data Augmentation DAIL-SQL [8] LLaMA* [11, 38] Aug-2023 VLDB’24
Symbol-LLM [83] CodeLLaMA [11] Nov-2023 ACL’24
CodeS [24] StarCoder [11, 1, 42] Feb-2024 SIGMOD’24
StructLM [84] CodeLLaMA [11] Feb-2024 COLM’24
Dubo-SQL [71] ChatGPT [1] Apr-2024 arXiv’24
Distillery [58] LLaMA-3.1* [1] Aug-2024 NeurIPS’24 Workshop
XiYan-SQL [85] Qwen2.5-Coder [11, 1] Nov-2024 arXiv’24
SHARE [86] LLaMA-3.1* [11, 1, 36, 38] Jun-2025 ACL’25
Multi-task Tuning SQL-LLaMA [48] CodeLLaMA [11, 1] Dec-2023 COLING’25
DTS-SQL [88] Mistral* [11, 37] Feb-2024 EMNLP’24 Findings
KaSLA [29] DeepSeek-Coder* [11, 1] Oct-2024 arXiv’24
MSc-SQL [89] Gemma-2 [11, 1] Oct-2024 NAACL’25
XiYan-SQL [85] Qwen2.5-Coder [11, 1] Nov-2024 arXiv’24
ROUTE [90] Qwen2.5* [11, 1, 37, 36, 38] Feb-2025 ICLR’25
SHARE [86] LLaMA-3.1* [11, 1, 36, 38] Jun-2025 ACL’25
Supervised Fine-tuning (SFT) is the widely-used approach within the FT paradigm for training LLMs [27, 113]. For open-source LLMs, like LLaMA [117] and Qwen [118] series, the most straightforward way to enable the model to adapt to a specific domain is to apply SFT over collected domain-specific data. The SFT phase typically serves as an initial step in a well-designed training pipeline [128, 129], including LLM-based text-to-SQL training. Consider the prompt defined in Eq. 4 and 5 as $P$, and let $Y=[y_{1},y_{2},\dots,y_{T}]$ represent the target ground-truth SQL query. The probability Pr for an LLM $\pi$ to generate $Y$ given $P$ is formally modeled as:
@@
As a vanilla fine-tuning approach for LLM-based text-to-SQL, SFT is widely adopted in text-to-SQL research across various open-source LLMs [24, 120, 8]. FT paradigms are often more foundational in LLM-based text-to-SQL than ICL, which is the mainstream approach in the community. This is largely due to the comparably limited capabilities of open-source LLMs versus proprietary models in understanding complex database content. For the existing well-designed methods in the FT paradigm, we group these methods into different categories based on their training mechanisms, as shown in Tab. IV.
@@ IV-B1 Enhanced Architecture
The widely-used Generative Pre-trained Transformer (GPT) framework [105] employs a decoder-only transformer architecture and conventional auto-regressive decoding for text generation. With advancements in the GPT framework, enhancements in LLM-based text-to-SQL involve modifications of the standard transformer backbone. These modifications incorporate tailored architecture to efficiently handle the structural and syntactic complexities of SQL. In traditional text-to-SQL research, this tailored architecture has effectively enhanced PLM-based performance. Techniques include schema-aware encoding [20], optimized attention [130], and targeted decoding [131], often relying on graph structures. Most recently, the speed of generating SQL queries by LLMs is notably slower than traditional PLM-based methods [19, 26], posing a challenge to building high-efficiency local NLIDBs.
As a solution, CLLMs [82] are designed to address this challenge with specialized decoding, offering speedup efficiency for SQL generation and reducing high latency.
@@ IV-B2 Pre-training
Pre-training is a fundamental phase of the complete training process, aimed at acquiring text generation capabilities through auto-regressive training on extensive data [132]. Code-specific LLMs (e.g., CodeLLaMA [87], StarCoder [133]) are pre-trained on code data, with diverse programming languages enabling them to generate code per user instructions [124, 134]. For SQL-specific pre-training, a key challenge is that SQL/database-related content makes up only a small part of the pre-training corpus. As a result, the open-source LLMs do not acquire a solid understanding of converting NL questions to SQL during their pre-training.
PLM-based methods propose initial attempts in pre-training paradigms, such as STAR [135] and multitask approaches [136], which improve SQL generation by leveraging SQL-guided and multitask pre-training strategies. For the most recent LLM-based text-to-SQL, the pre-training phase of the CodeS [24] model consists of three stages of incremental pre-training. Starting from a code-specific LLM [133], CodeS are incrementally pre-trained on a hybrid training corpus, including SQL-related data, NL-to-Code data, and NL-related data. The text-to-SQL understanding and performance are significantly improved.
@@ IV-B3 Data Augmentation
During the fine-tuning process, the quality of training labels is the most straightforward factor affecting the model’s performance [137]. Training with low-quality or absent labels is “making bricks without straw.” Using high-quality or augmented data for fine-tuning consistently outperforms the meticulously designed methods relying on low-quality or raw data [93, 27]. Data-augmented fine-tuning in text-to-SQL has made substantial progress by focusing on enhancing data quality and efficiency during the SFT process.
DAIL-SQL [8] is designed as an ICL framework that uses a sampling strategy to enhance few-shot instances. Incorporating these sampled instances in the SFT process boosts the performance of open-source LLMs. Symbol-LLM [83] proposes injection and infusion stages for data-augmented instruction tuning. CodeS [24] augments the training data with bi-directional generation aided by ChatGPT. StructLM [84] trains on multiple structured knowledge tasks to improve overall capability. Dubo-SQL [71] created concise prompt templates for fine-tuning, which reduces costs and simultaneously achieves record performance. Distillery [58] employs an iterative failure-aware approach, selecting new training samples based on error analysis. XiYan-SQL [85] uses basic-syntax SFT and generation-enhanced SFT, employing hybrid augmented training data to fine-tune the SQL generator. SHARE [86] designed a rule-based error perturbation to augment schema linking and final SQL inference based on the decomposed trajectories.

### v7

@@ IV-A3 $\mathbf{C}_{2}$-Prompt Optimization
1. Advanced few-shot sampling: This strategy optimizes example selection through semantic similarity, diversity, or domain alignment. DESEM [guo2023desem] and retrieval-augmented frameworks [guo2023ragsql, thorpe2024dubo, maamari2024the] de-semanticize questions to retrieve examples with matching intent skeletons or vector embeddings. DAIL-SQL [gao2023dailsql] balances quality and quantity by masking domain-specific tokens and ranking examples based on the Euclidean distance between questions and SQL queries, while SAFE-SQL [lee2025safe] combines skeleton masking with self-augmentation. ODIS [chang2023selective] hybridizes out-of-domain and synthetic in-domain examples for robustness, and FUSED [wang2024fused] enhances diversity through iterative demonstration synthesis. Dynamic selection is explored in ACT-SQL [zhang2023actsql] and MCS-SQL [lee2025mcs], which adaptively retrieve examples based on defined similarity scores or masked similarity scores.
2. Schema augmentation: These methods refine schema representations to reduce redundancy and improve relevance. C3 [dong2023c3] distills question-relevant schemas by removing unrelated tables or columns, while RSL-SQL [cao2024rsl] simplifies schemas using NL descriptions. ReFoRCE [deng2025reforce] employs table compression and column exploration to manage long-context schema complexity. Gen-SQL [shi2025gen] bridges NL-SQL gaps by generating pseudo-schemas that align with user questions. QDecomp [tai2023exploring] and SD+SA+Voting [nan2023enhancing] further augment prompts with schema-aware semantic and structural hints. For more complex databases, LinkAlign [wang2025linkalign] introduces a schema-based question rewriting strategy, combined with a retrieve-then-filter schema selection mechanism for accurate linking.
3. External knowledge integration: This group enhances prompts with domain-specific knowledge or feedback mechanisms. Knowledge-to-SQL [hong2024knowledgetosql] trains a DELLM to inject external knowledge about question-database relationships, while ICRL [toteja2025context] integrates reinforcement learning with schema-graph-derived knowledge bases. MAC-SQL [wang2024macsql] and DIN-SQL [pourreza2023dinsql] implicitly utilize decomposition patterns as procedural knowledge to guide generation. C3 [dong2023c3] and SAFE-SQL [lee2025safe] incorporate calibration bias and self-augmented feedback loops, respectively, to iteratively refine outputs.
@@ IV-A4 $\mathbf{C}_{3}$-Reasoning Enhancement
1. Chain-of-thought (CoT)-based reasoning: CoT-based methods guide LLMs in generating intermediate reasoning steps before predicting SQLs. While standard CoT prompts (e.g., “Let’s think step by step”) show limited effectiveness in text-to-SQL [gao2023dailsql, li2023BIRD], adaptations help bridge this gap. ACT-SQL [zhang2023actsql] automates CoT example generation by linking question slices to database columns via similarity matching. QDecomp [tai2023exploring] restructures SQL clauses into logical steps ordered by execution flow, reducing error propagation compared to generic CoT. CHASE-SQL [pourreza2025chasesql] extends this with multi-path reasoning for candidate generation using “divide and conquer” CoT strategy.
2. Consistency-driven reasoning: These methods improve reliability through candidate diversity and agreement metrics before execution. Self-consistency [wang2023selfconsistency] is employed in frameworks like C3 [dong2023c3], DAIL-SQL [gao2023dailsql], and SuperSQL [li2024dawn], where majority voting selects the most frequent SQL candidate based on syntax or structural similarity. SD+SA+Voting [nan2023enhancing] refines this by filtering candidates using schema-aware rules (e.g., valid joins, clause ordering) before voting. Unlike execution-dependent methods, these strategies prioritize logical consistency and syntactic validity over runtime feedback, making them lightweight but less robust to semantic errors.
3. Tool-augmented and agentic reasoning: Hybrid techniques combine SQL generation with external computation or tools. SQL-CRAFT [xia2024sqlcraft] employs Program-of-Thoughts [chen2023program] (PoT) prompting, requiring LLMs to generate Python code alongside SQL for arithmetic-heavy queries, improving performance on datasets like BIRD [li2023BIRD]. FUXI [gu2024fuxi] integrates dedicated tools for schema linking and syntax checking, dynamically invoking them during generation. These approaches differ from pure prompting strategies by explicitly bridging LLM reasoning with external symbolic systems. Also, recent frameworks such as ReFoRCE [deng2025reforce] and Spider-Agent [lei2025spider2] adopt step-by-step agentic reasoning, where LLMs act as specialized agents that iteratively plan actions to handle complex database scenarios.
@@ IV-A5 $\mathbf{C}_{4}$-Execution Refinement
1. Feedback-driven regeneration: These methods iteratively refine SQL queries using execution errors or intermediate results. Self-Debugging [chen2024selfdebugging] enables LLMs to autonomously correct errors via natural language explanations of failed executions, while DESEM [guo2023desem] uses error-type-specific prompts with termination criteria to prevent infinite loops. Multi-agent frameworks like MAC-SQL [wang2024macsql] and CHASE-SQL [pourreza2025chasesql] utilize agents (e.g., Refiner, Query Fixer) to diagnose SQLite errors and regenerate SQL fixes. SQL-CRAFT [xia2024sqlcraft] introduces adaptive control to balance correction depth, preventing over-refinement. Schema-focused methods like RSL-SQL [cao2024rsl] and DIN-SQL [pourreza2023dinsql] use generic or gentle prompts to help LLMs identify schema misalignments during regeneration. Knowledge-to-SQL [hong2024knowledgetosql] optimizes its DELLM model through preference learning from execution feedback, while FUXI [gu2024fuxi] integrates error feedback into tool-based reasoning. E-SQL [caferouglu2024esql] enhances questions with execution-aware candidate predictions. Agentic frameworks [lei2025spider2] also incorporate feedback-driven regeneration, where specialized agents iteratively analyze execution feedback to plan corrective actions to refine SQLs.
2. Execution-guided selection: This strategy leverages database execution results to select or rank candidates. ReFoRCE [deng2025reforce] and PET-SQL [li2024petsql] prune invalid SQL queries that fail execution, then apply bias-aware voting (e.g., difficulty-weighted or efficiency-prioritized as in MCS-SQL [lee2025mcs]). Probabilistic verification distinguishes methods: MRC-EXEC [shi2022natural] minimizes Bayes risk by comparing execution outputs to expected results, and LEVER [ni2023lever] trains a verifier to predict correctness likelihoods from execution traces. Hybrid approaches like the retrieval-augmented framework [guo2023ragsql] use execution feedback for iterative SQL revision through semantic gap analysis, bridging generation and selection.
@@ IV-B Fine-tuning
Enhanced Architecture CLLMs [kou2024cllms] DeepSeek* [yu2018Spider] Mar-2024 ICML’24
Pre-training CodeS [li2024codes] StarCoder [yu2018Spider, li2023BIRD] Feb-2024 SIGMOD’24
Data Augmentation DAIL-SQL [gao2023dailsql] LLaMA* [yu2018Spider, deng2021Spider-Realistic] Aug-2023 VLDB’24
Symbol-LLM [xu2024symbolllm] CodeLLaMA [yu2018Spider] Nov-2023 ACL’24
CodeS [li2024codes] StarCoder [yu2018Spider, li2023BIRD, lei2025spider2] Feb-2024 SIGMOD’24
StructLM [zhuang2024structlm] CodeLLaMA [yu2018Spider] Feb-2024 COLM’24
Dubo-SQL [thorpe2024dubo] ChatGPT [li2023BIRD] Apr-2024 arXiv’24
Distillery [maamari2024the] LLaMA-3.1* [li2023BIRD] Aug-2024 NeurIPS’24 Workshop
XiYan-SQL [gao2024xiyan] Qwen2.5-Coder [yu2018Spider, li2023BIRD] Nov-2024 arXiv’24
SHARE [qu2025share] LLaMA-3.1* [yu2018Spider, li2023BIRD, gan2021Spider-DK, deng2021Spider-Realistic] Jun-2025 ACL’25
Multi-task Tuning SQL-LLaMA [wang2024macsql] CodeLLaMA [yu2018Spider, li2023BIRD] Dec-2023 COLING’25
DTS-SQL [pourreza2024dtssql] Mistral* [yu2018Spider, gan2021Spider-SYN] Feb-2024 EMNLP’24 Findings
KaSLA [yuan2025knapsack] DeepSeek-Coder* [yu2018Spider, li2023BIRD] Oct-2024 arXiv’24
MSc-SQL [gorti2024mscsql] Gemma-2 [yu2018Spider, li2023BIRD] Oct-2024 NAACL’25
XiYan-SQL [gao2024xiyan] Qwen2.5-Coder [yu2018Spider, li2023BIRD] Nov-2024 arXiv’24
ROUTE [qin2025route] Qwen2.5* [yu2018Spider, li2023BIRD, gan2021Spider-SYN, gan2021Spider-DK, deng2021Spider-Realistic] Feb-2025 ICLR’25
SHARE [qu2025share] LLaMA-3.1* [yu2018Spider, li2023BIRD, gan2021Spider-DK, deng2021Spider-Realistic] Jun-2025 ACL’25
Supervised Fine-tuning (SFT) is the widely-used approach within the FT paradigm for training LLMs [zhao2023survey, zheng2024llamafactory]. For open-source LLMs, like LLaMA [touvron2023llama2] and Qwen [bai2023qwen] series, the most straightforward way to enable the model to adapt to a specific domain is to apply SFT over collected domain-specific data. The SFT phase typically serves as an initial step in a well-designed training pipeline [rafailo2023direct, ouyang2022training], including LLM-based text-to-SQL training. Consider the prompt defined in Eq. 4 and 5 as $P$, and let $Y=[y_{1},y_{2},\dots,y_{T}]$ represent the target ground-truth SQL query. The probability Pr for an LLM $\pi$ to generate $Y$ given $P$ is formally modeled as:
@@
As a vanilla fine-tuning approach for LLM-based text-to-SQL, SFT is widely adopted in text-to-SQL research across various open-source LLMs [li2024codes, zhang2024benchmarking, gao2023dailsql]. FT paradigms are often more foundational in LLM-based text-to-SQL than ICL, which is the mainstream approach in the community. This is largely due to the comparably limited capabilities of open-source LLMs versus proprietary models in understanding complex database content. For the existing well-designed methods in the FT paradigm, we group these methods into different categories based on their training mechanisms, as shown in Tab. IV.
@@ IV-B1 Enhanced Architecture
The widely-used Generative Pre-trained Transformer (GPT) framework [radford2018improving] employs a decoder-only transformer architecture and conventional auto-regressive decoding for text generation. With advancements in the GPT framework, enhancements in LLM-based text-to-SQL involve modifications of the standard transformer backbone. These modifications incorporate tailored architecture to efficiently handle the structural and syntactic complexities of SQL. In traditional text-to-SQL research, this tailored architecture has effectively enhanced PLM-based performance. Techniques include schema-aware encoding [li2023graphix], optimized attention [gong2024graph], and targeted decoding [cao2023heterogeneous], often relying on graph structures. Most recently, the speed of generating SQL queries by LLMs is notably slower than traditional PLM-based methods [li2023resdsql, yang2024harnessing], posing a challenge to building high-efficiency local NLIDBs.
As a solution, CLLMs [kou2024cllms] are designed to address this challenge with specialized decoding, offering speedup efficiency for SQL generation and reducing high latency.
@@ IV-B2 Pre-training
Pre-training is a fundamental phase of the complete training process, aimed at acquiring text generation capabilities through auto-regressive training on extensive data [erhan2010does]. Code-specific LLMs (e.g., CodeLLaMA [roziere2023codellama], StarCoder [li2023starcoder]) are pre-trained on code data, with diverse programming languages enabling them to generate code per user instructions [chen2021evaluating, yufei2023aligning]. For SQL-specific pre-training, a key challenge is that SQL/database-related content makes up only a small part of the pre-training corpus. As a result, the open-source LLMs do not acquire a solid understanding of converting NL questions to SQL during their pre-training.
PLM-based methods propose initial attempts in pre-training paradigms, such as STAR [cai2022star] and multitask approaches [giaquinto2023multitask], which improve SQL generation by leveraging SQL-guided and multitask pre-training strategies. For the most recent LLM-based text-to-SQL, the pre-training phase of the CodeS [li2024codes] model consists of three stages of incremental pre-training. Starting from a code-specific LLM [li2023starcoder], CodeS are incrementally pre-trained on a hybrid training corpus, including SQL-related data, NL-to-Code data, and NL-related data. The text-to-SQL understanding and performance are significantly improved.
@@ IV-B3 Data Augmentation
During the fine-tuning process, the quality of training labels is the most straightforward factor affecting the model’s performance [song2022learningnoise]. Training with low-quality or absent labels is “making bricks without straw.” Using high-quality or augmented data for fine-tuning consistently outperforms the meticulously designed methods relying on low-quality or raw data [deng2022recent, zhao2023survey]. Data-augmented fine-tuning in text-to-SQL has made substantial progress by focusing on enhancing data quality and efficiency during the SFT process.
DAIL-SQL [gao2023dailsql] is designed as an ICL framework that uses a sampling strategy to enhance few-shot instances. Incorporating these sampled instances in the SFT process boosts the performance of open-source LLMs. Symbol-LLM [xu2024symbolllm] proposes injection and infusion stages for data-augmented instruction tuning. CodeS [li2024codes] augments the training data with bi-directional generation aided by ChatGPT. StructLM [zhuang2024structlm] trains on multiple structured knowledge tasks to improve overall capability. Dubo-SQL [thorpe2024dubo] created concise prompt templates for fine-tuning, which reduces costs and simultaneously achieves record performance. Distillery [maamari2024the] employs an iterative failure-aware approach, selecting new training samples based on error analysis. XiYan-SQL [gao2024xiyan] uses basic-syntax SFT and generation-enhanced SFT, employing hybrid augmented training data to fine-tune the SQL generator. SHARE [qu2025share] designed a rule-based error perturbation to augment schema linking and final SQL inference based on the decomposed trajectories.